Resolve the 29-lb force F into components along the u and v axes. 0 15% 30° lu
Part A Determine the magnitude of u component. Express your answer to three significant figures and include the appropriate units.
Part B Determine the magnitude of v component. Express your answer to three significant figures and include the appropriate units.
Answer #1

3 15 1y S 6 lu

Part A)

In triangle , ACD

AD = Fu = component of the force along u-axis

AC = force F = 29 lb

using law of Sines

AC/Sin105 = AD/Sin45

F/Sin105 = Fu/Sin45

29/Sin105 = Fu/Sin45

Fu = 21.2 lb

Part B)

In triangle , ABC

AB = Fv = component of the force along v-axis

AC = force F = 29 lb

using law of Sines

AC/Sin105 = AB/Sin30

F/Sin105 = Fv/Sin30

29/Sin105 = Fv/Sin30

Fv = 15.0 lb